# Vevo2: A Unified and Controllable Framework for Speech and Singing Voice Generation

### v2

@@ V-C Controllability over Timbre and Style
Zero-Shot SVC For zero-shot SVC, we include five baselines: three models pre-trained only on speech data (FACodec [naturalspeech3], Vevo [vevo], and CosyVoice 2 [cosyvoice2]), and two models pre-trained on both speech and singing voice data (NeuCoSVC 2 [neucosvc] and SeedVC (SVC) [seedvc]). Notably, for Vevo and CosyVoice 2, we utilize their flow-matching (FM) models. Similarly, we include Vevo2’s FM model as a baseline (denoted as Vevo2-FM). The results in Table III reveal two key findings:
@@
(2) Comparing Vevo2-FM and Vevo2, there are two interesting observations: a) Vevo2 shows significant improvements in intelligibility, which we attribute to both the explicit use of text input and intelligibility alignment during post-training. b) Although Vevo2’s Melody-MOS is lower than Vevo2-FM’s, its absolute value still exceeds 2, indicating its ability to “roughly follow the melody contour”. This trade-off benefits style-converted SVC [^12] For style-converted SVC, we aim for the generated singing voice to follow the source’s coarse-grained melody contour while imitating fine-grained expression details (such as vibrato) from the reference [vevo, svcc2025].: specifically, Vevo2 achieves better style imitation performance (higher Style-CMOS), leading to improved speaker similarity (higher SIM). Similar characteristics are also observed in style-converted VC tasks [vevo].
Zero-Shot Accent and Emotion Conversion Table V demonstrates Vevo2’s performance on zero-shot accent and emotion conversion tasks. We benchmarked against Vevo [vevo] and adopted their evaluation sets and two style conversion metrics: Accent-SIM (computed using a pre-trained accent classifier [^13] https://huggingface.co/Jzuluaga/accent-id-commonaccent_ecapa) and Emotion-SIM (calculated via a pre-trained emotion classifier [^14] https://github.com/ddlBoJack/emotion2vec) to assess style conversion performance. The results indicate that Vevo2 achieves comparable style similarity scores to Vevo while exhibiting superior performance in both intelligibility and speaker similarity metrics.
@@
Figure 5: Effect of intelligibility reward (w/ Intell) and prosody similarity reward (w/ Prosody) for post-training. The right figure presents subjective evaluation results on the instrument-to-sing task.
V-D Controllability over Duration
In our experiments, we investigated Vevo2’s duration control effectiveness across three tasks: speech editing (SE), singing lyric editing (SLE), and style-converted SVC. We adopt the metric, Differences of Duration (DDUR, $\downarrow$) [non-parallel-seq2seq-vc, vevo], which measures the average absolute difference in seconds between the actual generated duration ($\hat{d}$) and the target duration ($d$). Additionally, we introduce Duration Consistency (DC, $\uparrow$) to evaluate the relative accuracy of duration control:
For editing tasks, following F5-TTS [f5tts], we calculate the target duration based on the linear relationship between the lengths of edited and raw text. For the style-converted SVC, we constrain the target duration to match the source waveform duration. As demonstrated in Table V, Vevo2 achieved over 97% duration consistency across all three tasks, exhibiting good duration controllability despite its autoregressive nature.
@@ V-E Controllability over Pitch Region
The pitch region control is particularly effective for voice conversion (VC) and singing voice conversion (SVC) tasks [amphion-svc, svcc-vits-ziqian, svcc-2023]. For Vevo2, to achieve the pitch region control:
@@ V-F Effectiveness of Multi-Objective Post-Training
The evaluation sets for humming-to-singing and instrument-to-singing tasks are constructed through the following procedure. For melody sources, we selected humming samples [humtrans] and instrumental sounds, including flute [flute-dataset] and violin [violin-dataset] recordings. For each melody sample, we leveraged its MIDI annotation’s note count to generate appropriate lyrics using DeepSeek-v3. The following prompt was used for DeepSeek-v3: "I will provide the number of notes in a melody. Please write suitable lyrics for it. Note: (1) The number of words in the lyrics should roughly match the number of notes; (2) Please only return the result, without any further explanation." After obtaining both melody and lyrics, we randomly assigned a target singer from SingStyle111 [singstyle111] for each sample, thereby creating evaluation triplets in the form of (melody, lyric, singer). The evaluation set comprises 1,000 samples in total, evenly distributed between English (500) and Chinese (500) lyrics.
@@ Appendix A Details of Baselines
  • F5-TTS [f5tts]: It adopts a single-stage flow-matching transformer to predict Mel spectrograms from text. It is pre-trained on Emilia [emilia] (101K hours of speech data). We use the officially released checkpoint [^15] https://huggingface.co/SWivid/F5-TTS to generate samples.
  • MaskGCT [maskgct]: It adopts a two-stage pipeline: text to content-style tokens (50 Hz), and content-style tokens to acoustic tokens. It is pre-trained on Emilia [emilia] (101K hours of speech data). We use the officially released checkpoint [^16] https://huggingface.co/amphion/MaskGCT to generate samples.
  • CosyVoice 2 [cosyvoice]: It contains an AR transformer to predict the content-style tokens (25 Hz) from text, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^17] https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B to generate samples, which is pre-trained on 167K hours of private data.
@@
  • StyleSinger [stylesinger]: It is a zero-shot SVS model for style transfer that is pre-trained on the Chinese GTSinger data [gtsinger]. We use the officially released checkpoint [^18] https://github.com/AaronZ345/StyleSinger to generate samples.
  • TCSinger [tcsinger]: It is a zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. It is pre-trained on the GTSinger data [gtsinger]. We use the officially released checkpoint [^19] https://github.com/AaronZ345/TCSinger to generate samples.
@@
Model
#Hz
Data
(#hours)
PT?
Regular Speech
WER SIM UTMOS
Ground Truth - - - 1.699 0.740 3.158
F5-TTS [f5tts] - 101K 3.442 0.670 3.090
MaskGCT [maskgct] 50 101K 2.344 0.738 3.136
CosyVoice2 [cosyvoice] 25 167K 2.091 0.709 3.685
Vevo2-base
12.5
101K 6.479 0.719 3.378
7K 29.572 0.654 3.081
101K, 7K 5.201 0.720 3.386
Vevo2 12.5 101K, 7K 3.292 0.724 3.405
@@
  • FACodec [naturalspeech3]: It adopts an auto-encoder and residual vector quantization based architecture. It decouples the raw waveform into factorized attributes through ASR, F0 prediction, and speaker classification tasks, trained on the Libri-light dataset [libri-light]. We use the released checkpoint in Amphion  [^23] https://huggingface.co/amphion/naturalspeech3_facodec [amphion, amphion_v0.2] (which is implemented by the authors) to generate samples.
  • Vevo-FM [vevo]: It is the flow-matching model of Vevo (i.e., Vevo-Timbre in the original paper), which uses a 50 Hz VQ-VAE content-style tokenizer based on HuBERT [hubert] representations. It is pre-trained on Emilia (101K hours of speech data). We use the officially released checkpoint [^24] https://huggingface.co/amphion/Vevo to generate samples.
  • CosyVoice2-FM [cosyvoice2]: It is the flow-matching model of CosyVoice 2 that uses 25 Hz content-style tokenizer, consistent with the configuration described in the zero-shot TTS task above.
  • NeuCoSVC 2 [neucosvc]: It is an enhanced version of NeuCoSVC and is pre-trained on an extensive internal dataset comprising approximately 500 hours of singing voice data and various open-source speech datasets [neucosvc]. We use the officially released checkpoint [^25] https://github.com/thuhcsi/NeuCoSVC/tree/NeuCoSVC2 to generate samples.
  • SeedVC [seedvc]: It is a diffusion transformer that uses an external timbre shifter during training to perturb the source speech timbre. We use the officially released pre-trained VC and SVC models [^26] https://github.com/Plachtaa/seed-vc, which are denoted as SeedVC (VC) and SeedVC (SVC), to conduct the inference.
@@
  • Vevo [vevo]: It is named as Vevo-Style in the original paper, which contains an AR transformer to predict content-style tokens (50 Hz) from content tokens, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^27] https://huggingface.co/amphion/Vevo to generate samples, which is pre-trained on 101K hours of Emilia.
@@ Appendix B Additional Experimental Results
B-A Zero-Shot Text to Speech
@@
The zero-shot TTS evaluation results for expressive speech and singing voice domains are presented in Table I. Table VII demonstrates the performance on the regular speech domain using the SeedTTS evaluation sets [seedtts], a widely adopted benchmark in the TTS field [maskgct, cosyvoice2, f5tts]. The results corroborate our findings from Section V-A. Specifically, we observe that incorporating singing voice data alongside speech data enhances performance across all metrics (WER, SIM, UTMOS) in the regular speech domain. Notably, the post-trained Vevo2 achieves comparable performance to existing state-of-the-art baselines, despite utilizing a content-style tokenizer with a lower frame rate.
B-B Zero-Shot Voice Conversion
Table VIII presents the results of our voice conversion evaluation, utilizing the same baselines [^28] For SeedVC, we employed the official SVC checkpoint for the zero-shot SVC task (Table III) and the official VC checkpoint for the zero-shot VC task (Table VIII). as the SVC task in Table III, where we use the SeedTTS VC evaluation set [seedtts]. The VC task results lead to conclusions similar to those observed in the SVC experiments: First, Vevo2-FM demonstrates robust timbre disentanglement capabilities, with speaker similarity (SIM) metrics matching or exceeding existing VC models, particularly for English content. Furthermore, the transition from Vevo2-FM to Vevo2 (i.e., from style-preserved VC to style-converted VC) yields additional improvements in speaker similarity. Moreover, the explicit incorporation of text input significantly enhances intelligibility performance.
@@ B-B Zero-Shot Voice Conversion
FACodec [naturalspeech3] 4.88 0.355 2.79 8.20 0.495 2.10
Vevo-FM [vevo] 3.01 0.536 3.51 4.06 0.695 2.86
CosyVoice2-FM [cosyvoice2] 4.66 0.530 3.86 2.80 0.728 3.08
NeuCoSVC 2 [neucosvc] 3.57 0.227 3.15 2.90 0.468 2.51
SeedVC (VC) [seedvc] 2.97 0.565 3.31 2.45 0.737 2.69

### v3

@@ V-C Controllability over Timbre and Style
Zero-Shot SVC For zero-shot SVC, we include five baselines: three models pre-trained only on speech data (FACodec [36], Vevo [95], and CosyVoice 2 [22]), and two models pre-trained on both speech and singing voice data (NeuCoSVC 2 [59] and SeedVC (SVC) [46]). Notably, for Vevo and CosyVoice 2, we utilize their flow-matching (FM) models. Similarly, we include Vevo2’s FM model as a baseline (denoted as Vevo2-FM). The results in Table III reveal two key findings:
@@
(2) Comparing Vevo2-FM and Vevo2, there are two interesting observations: a) Vevo2 shows significant improvements in intelligibility, which we attribute to both the explicit use of text input and intelligibility alignment during post-training. b) Although Vevo2’s Melody-MOS is lower than Vevo2-FM’s, its absolute value still exceeds 2, indicating its ability to “roughly follow the melody contour”. This trade-off benefits style-converted SVC [^12] For style-converted SVC, we aim for the generated singing voice to follow the source’s coarse-grained melody contour while imitating fine-grained expression details (such as vibrato) from the reference [95, 68].: specifically, Vevo2 achieves better style imitation performance (higher Style-CMOS), leading to improved speaker similarity (higher SIM). Similar characteristics are also observed in style-converted VC tasks [95].
Zero-Shot Accent and Emotion Conversion Table V demonstrates Vevo2’s performance on zero-shot accent and emotion conversion tasks. We benchmarked against Vevo [95] and adopted their evaluation sets and two style conversion metrics: Accent-SIM (computed using a pre-trained accent classifier [^13] https://huggingface.co/Jzuluaga/accent-id-commonaccent_ecapa) and Emotion-SIM (calculated via a pre-trained emotion classifier [^14] https://github.com/ddlBoJack/emotion2vec) to assess style conversion performance. The results indicate that Vevo2 achieves comparable style similarity scores to Vevo while exhibiting superior performance in both intelligibility and speaker similarity metrics.
V-D Controllability over Duration
In our experiments, we investigated Vevo2’s duration control effectiveness across three tasks: speech editing (SE), singing lyric editing (SLE), and style-converted SVC. We adopt the metric, Differences of Duration (DDUR, $\downarrow$) [87, 95], which measures the average absolute difference in seconds between the actual generated duration ($\hat{d}$) and the target duration ($d$). Additionally, we introduce Duration Consistency (DC, $\uparrow$) to evaluate the relative accuracy of duration control:
For editing tasks, following F5-TTS [11], we calculate the target duration based on the linear relationship between the lengths of edited and raw text. For the style-converted SVC, we constrain the target duration to match the source waveform duration. As demonstrated in Table V, Vevo2 achieved over 97% duration consistency across all three tasks, exhibiting good duration controllability despite its autoregressive nature.
@@ V-D Controllability over Duration
Figure 5: Effect of intelligibility reward (w/ Intell) and prosody similarity reward (w/ Prosody) for post-training. The right figure presents subjective evaluation results on the instrument-to-singing task.
@@ V-E Controllability over Pitch Region
The pitch region control is particularly effective for voice conversion (VC) and singing voice conversion (SVC) tasks [91, 52, 33]. For Vevo2, to achieve the pitch region control:
@@ V-F Effectiveness of Multi-Objective Post-Training
The evaluation sets for humming-to-singing and instrument-to-singing tasks are constructed through the following procedure. For melody sources, we selected humming samples [45] and instrumental sounds, including flute [8] and violin [20] recordings. For each melody sample, we leveraged its MIDI annotation’s note count to generate appropriate lyrics using DeepSeek-v3. The following prompt was used for DeepSeek-v3: "I will provide the number of notes in a melody. Please write suitable lyrics for it. Note: (1) The number of words in the lyrics should roughly match the number of notes; (2) Please only return the result, without any further explanation." After obtaining both melody and lyrics, we randomly assigned a target singer from SingStyle111 [16] for each sample, thereby creating evaluation triplets in the form of (melody, lyric, singer). The evaluation set comprises 1,000 samples in total, evenly distributed between English (500) and Chinese (500) lyrics.
@@ VI Conclusion
Acknowledgment
This work is partially supported by the National Natural Science Foundation of China under Grant 62376237, 2023 Shenzhen Stability Science Program, Internal Project Fund from Shenzhen Research Institute of Big Data (Grant No. T00120230002), and the Program for Guangdong Introducing Innovative and Enterpreneurial Teams (Grant No. 2023ZT10X044). We thank Yicheng Gu, Yushun Zhang and the anonymous reviewers for their insightful comments and suggestions. We appreciate the efforts of all the subjects during the subjective evaluation.
@@ Appendix A Details of Baselines
  • F5-TTS [11]: It adopts a single-stage flow-matching transformer to predict Mel spectrograms from text. It is pre-trained on Emilia [30] (101K hours of speech data). We use the officially released checkpoint [^15] https://huggingface.co/SWivid/F5-TTS to generate samples.
  • MaskGCT [71]: It adopts a two-stage pipeline: text to content-style tokens (50 Hz), and content-style tokens to acoustic tokens. It is pre-trained on Emilia [30] (101K hours of speech data). We use the officially released checkpoint [^16] https://huggingface.co/amphion/MaskGCT to generate samples.
  • CosyVoice 2 [21]: It contains an AR transformer to predict the content-style tokens (25 Hz) from text, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^17] https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B to generate samples, which is pre-trained on 167K hours of private data.
@@
  • StyleSinger [97]: It is a zero-shot SVS model for style transfer that is pre-trained on the Chinese GTSinger data [99]. We use the officially released checkpoint [^18] https://github.com/AaronZ345/StyleSinger to generate samples.
  • TCSinger [98]: It is a zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. It is pre-trained on the GTSinger data [99]. We use the officially released checkpoint [^19] https://github.com/AaronZ345/TCSinger to generate samples.
@@
Model #Hz Data (#hours) PT? SeedTTS test-en SeedTTS test-zh
WER SIM UTMOS WER SIM UTMOS
Ground Truth - - - 2.143 0.730 3.523 1.254 0.750 2.793
F5-TTS [11] - 101K 3.015 0.629 3.508 3.868 0.710 2.671
MaskGCT [71] 50 101K 2.403 0.711 3.590 2.284 0.765 2.682
CosyVoice2 [21] 25 167K 2.890 0.660 4.100 1.292 0.757 3.270
Vevo2-base 12.5 101K 5.846 0.687 3.749 7.112 0.752 3.007
7K 42.150 0.581 3.260 16.994 0.726 2.902
101K, 7K 5.770 0.686 3.763 4.632 0.752 3.009
Vevo2 12.5 101K, 7K 3.639 0.693 3.814 2.944 0.754 2.997
@@
  • FACodec [36]: It adopts an auto-encoder and residual vector quantization based architecture. It decouples the raw waveform into factorized attributes through ASR, F0 prediction, and speaker classification tasks, trained on the Libri-light dataset [37]. We use the released checkpoint in Amphion [^23] https://huggingface.co/amphion/naturalspeech3_facodec [94, 42] (which is implemented by the authors) to generate samples.
  • Vevo-FM [95]: It is the flow-matching model of Vevo (i.e., Vevo-Timbre in the original paper), which uses a 50 Hz VQ-VAE content-style tokenizer based on HuBERT [32] representations. It is pre-trained on Emilia (101K hours of speech data). We use the officially released checkpoint [^24] https://huggingface.co/amphion/Vevo to generate samples.
  • CosyVoice2-FM [22]: It is the flow-matching model of CosyVoice 2 that uses 25 Hz content-style tokenizer, consistent with the configuration described in the zero-shot TTS task above.
  • NeuCoSVC 2 [59]: It is an enhanced version of NeuCoSVC and is pre-trained on an extensive internal dataset comprising approximately 500 hours of singing voice data and various open-source speech datasets [59]. We use the officially released checkpoint [^25] https://github.com/thuhcsi/NeuCoSVC/tree/NeuCoSVC2 to generate samples.
  • SeedVC [46]: It is a diffusion transformer that uses an external timbre shifter during training to perturb the source speech timbre. We use the officially released pre-trained VC and SVC models [^26] https://github.com/Plachtaa/seed-vc, which are denoted as SeedVC (VC) and SeedVC (SVC), to conduct the inference.
@@
  • Vevo [95]: It is named as Vevo-Style in the original paper, which contains an AR transformer to predict content-style tokens (50 Hz) from content tokens, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^27] https://huggingface.co/amphion/Vevo to generate samples, which is pre-trained on 101K hours of Emilia.
@@ Appendix B Additional Experimental Results
B-A Visualization of Different Representations for Modeling Timbre and Melody
In Section III-A, we discretize chromagram features into prosody tokens via a VQ-VAE tokenizer. This design is motivated by two considerations. First, discrete tokens naturally fit the sequence modeling paradigm of auto-regressive (AR) transformers, enabling unified conditioning with text and other discrete representations. Second, we hypothesize that vector quantization can serve as an information bottleneck [57, 95, 88]: it encourages the model to compress salient melodic/prosodic cues into a compact discrete space, while suppressing residual acoustic characteristics (e.g., timbre or instrument-dependent patterns) that may still be present in continuous chromagram features.
To examine the above hypothesis, we conduct a controlled visualization experiment based on M4Singer [89]. We randomly select a set of MIDI annotations (i.e., melodies) and render each MIDI into multiple instrumental audio tracks using pretty_midi [^28] https://github.com/craffel/pretty-midi with different General MIDI instrument programs. We also include the original vocal utterance when available. This construction yields a controlled collection where the melody is fixed while the timbre varies across instruments.
For each audio sample in this collection, we extract and compare multiple representations: (1) the Mel spectrogram, (2) the chromagram features, and (3) our VQ-quantized chromagram tokens produced by the proposed prosody tokenizer. To obtain an utterance-level embedding for visualization, we apply mean pooling over time for Mel spectrogram and chromagram features. For VQ tokens, we first map each token index to its corresponding codebook embedding vector, and then mean-pool the resulting embedding sequence over time. Finally, we apply PCA to project all utterance-level embeddings into 2D for visualization. The results are shown in Figure 6 and 7.
(a) Mel Spectrogram
(b) Chromagram
(c) Prosody Token
Figure 6: Timbre invariance of different representations.
(a) Mel Spectrogram
(b) Chromagram
(c) Prosody Token
Figure 7: Melodic clusterability of different representations.
We highlight two properties that are desirable for a prosody/melody representation in controllable voice generation:
(1) Timbre invariance. A robust prosody representation should be insensitive to timbre so that it generalizes across speakers and instruments. In Figure 6, the Mel spectrogram (Figure 6(a)) exhibits noticeable clustering by instrument type, indicating strong timbre dependence. In contrast, the dense chromagram (Figure 6(b)) already mitigates this bias by substantially mixing samples from different instruments. Notably, the VQ-quantized chromagram, i.e., our proposed prosody token (Figure 6(c)), further increases the mixing across instruments, suggesting that discretization helps reduce residual timbre cues and promotes a more timbre-invariant representation.
(2) Melodic clusterability. An effective prosody representation should map identical melodies to compact regions in the embedding space regardless of timbre variations. Figure 7 shows that Mel spectrogram (Figure 7(a)) fail to form coherent clusters by melody ID, likely due to their entanglement with diverse acoustic factors. Dense chromagram features (Figure 7(b)) yield clearer melodic grouping, indicating improved melody sensitivity. Moreover, the VQ-quantized chromagram, i.e., our proposed prosody token (Figure 7(c)), produces more compact and better-separated clusters, implying that quantization preserves melodic structure while providing a robust discrete conditioning signal for controllable generation.
B-B Zero-Shot Text to Speech
@@
The zero-shot TTS evaluation results for expressive speech and singing voice domains are presented in Table I. Table VII demonstrates the performance on the regular speech domain using the SeedTTS evaluation sets [1], a widely adopted benchmark in the TTS field [71, 22, 11]. The results corroborate our findings from Section V-A. Specifically, we observe that incorporating singing voice data alongside speech data enhances performance across all metrics (WER, SIM, UTMOS) in the regular speech domain. Notably, the post-trained Vevo2 achieves comparable performance to existing state-of-the-art baselines, despite utilizing a content-style tokenizer with a lower frame rate.
B-C Zero-Shot Voice Conversion
@@
FACodec [36] 4.88 0.355 2.79 8.20 0.495 2.10
Vevo-FM [95] 3.01 0.536 3.51 4.06 0.695 2.86
CosyVoice2-FM [22] 4.66 0.530 3.86 2.80 0.728 3.08
NeuCoSVC 2 [59] 3.57 0.227 3.15 2.90 0.468 2.51
SeedVC (VC) [46] 2.97 0.565 3.31 2.45 0.737 2.69
@@
Table VIII presents the results of our voice conversion evaluation, utilizing the same baselines [^29] For SeedVC, we employed the official SVC checkpoint for the zero-shot SVC task (Table III) and the official VC checkpoint for the zero-shot VC task (Table VIII). as the SVC task in Table III, where we use the SeedTTS VC evaluation set [1]. The VC task results lead to conclusions similar to those observed in the SVC experiments: First, Vevo2-FM demonstrates robust timbre disentanglement capabilities, with speaker similarity (SIM) metrics matching or exceeding existing VC models, particularly for English content. Furthermore, the transition from Vevo2-FM to Vevo2 (i.e., from style-preserved VC to style-converted VC) yields additional improvements in speaker similarity. Moreover, the explicit incorporation of text input significantly enhances intelligibility performance.